# Temperature Dependent Mechanical and Structural Properties of Uniaxially Strained Planar Graphene

### v1

@@ 3.1 Temperature calibration
After inserting an energy density $e_{N}$ in the strained lattice, we observe that initially $T$ increases with time from its zero starting value and then following some relatively large fluctuations the system is settled at thermal equilibrium after at most $10^{3}~{}t_{u}$. The temperature $T(t)$ as a function of time is computed in our microcanonical MD simulations through the energy equipartition relation
@@
In order to test that thermal equilibrium has been reached, we compare the mean and standard deviation of the fluctuating temperature over various time windows. Before achieving thermal equilibrium the standard deviation of the time-averaged $T$ is relatively large and also changes depending on the time window. When thermal equilibrium is reached the temperature fluctuations and the standard deviation are consistently small. The mean temperature at thermal equilibrium is calculated by averaging over both the individual realizations and the recording time-window. We denote this average value $\overline{\left\langle T\right\rangle}$ by $T_{ave}$. In this case, the standard deviation of the measured values is computed using all data points over realizations and time.
The relationship between the additional energy density $e_{N}$ above the relaxed equilibrium loaded structures and the averaged temperature $T_{ave}$ is linear for all cases examined here, corresponding to temperatures up to 700 K. One representative case is shown in Figure 2. The resulting slopes of the data fittings are very close for all values of stress $\sigma$ (a difference in the computed values is observed only in the fourth significant digit) and they are slightly above $2k_{B}$ due to the nonlinearities of the potential energy. For finite loads, the slope slightly increases with the amount of stress. For a given value of stress $\sigma$, we use the slope of the $e_{N}$ versus $T_{ave}$ fitting in order to control temperature (within a 1% accuracy) in our investigation. In particular, we are setting the amount of the added energy density $e_{N}$ according to the desired temperature value.
Figure 2: Relation between the average temperature $T_{ave}$ at thermal equilibrium, evaluated through averaging over both time and the different realizations, and the energy density $e_{N}$ above the relaxed $T=0$ K graphene structure subjected to uniaxial tensile stress $\sigma_{z}=2.16$ eV/Å2 along the zigzag direction. One standard deviation of the $T_{ave}$ measurements is indicated by blue error-bars. The slope of the linear fitting of the presented data points (indicated by the gray line) equals to $1.74\times 10^{-4}$ eV/K.
In order to avoid potential edge effects and represent the behavior of bulk graphene, we collect data from the central region of the lattice to investigate the system’s elastic and structural properties discussed below. In particular, this sub-lattice has an analogous geometry to the larger structure, with number of columns $n_{I}=44$ and number of rows $n_{J}=45$.
@@ 3.2 Mechanical response
In the stress-controlled numerical implementation used here to examine the mechanical response of graphene, we need to compute the resulting strain due to the fixed force applied at the appropriate edge atoms. The uniaxial strain $\epsilon$ is obtained through the strain of the central row of the graphene sheet when the stress is applied along the armchair direction, while it is calculated by the average strain on the two central columns of the sheet in case of stress along the zigzag direction (see Figure 1). By $\epsilon_{T}$ is indicated the strain corresponding to temperature $T$.
@@
In the zero temperature case, $T=0$ K, the uniaxial strain $\epsilon_{0}$ in the bulk is determined through the relative change of the length of the central row (two central columns) of graphene subjected to the applied stress along the armchair (zigzag) direction, with respect to the length of the central row (columns) at the unstrained equilibrium configuration shown in Figure 1, where we take these measurements in the central region of the lattice. We clarify that for any length computations discussed here, the horizontal (vertical) length is measured as the difference of the $x$ ($y$) coordinates of the considered edge atoms. The stress-strain response is obtained in this way at 0 K
@@
When the temperature of the system is raised at finite values by adding energy to the equilibrated graphene, the lattice stretches and compresses in an oscillatory manner. The details of these oscillations depend on the temperature and the applied stress and will be discussed in a future publication. In this case one has to take into account that the strain measurement $\epsilon_{T}(t)$ is now exhibiting temporal oscillations. Since we consider 10-20 realizations of the random initial energy distribution, we register the average, over these realizations, strain in time $\left\langle\epsilon_{T}(t)\right\rangle$, noting that the aforementioned oscillations are in-phase in the different realizations. For evaluating the strain $\epsilon_{T}(t)$, the reference length $\ell_{T}^{r}$ corresponding to zero stress $\sigma=0$ at temperature $T$ is obtained by calculating the average, over realizations and time, of the length of the central row or columns of the sheet in the absence of any load. Then, when a stress is applied the time evolution of strain in a particular realization is computed as
$\epsilon_{T}(t)=\frac{\ell(t)-\ell_{T}^{r}}{\ell_{T}^{r}}.$ (6)
@@
where $\ell(t)$ is the length in time of the central row or the average length of the two central columns depending on the direction of the applied uniaxial load.
@@
In Figure 3 we highlight the behavior of $\left\langle{\epsilon_{T}}(t)\right\rangle$ for various values of stress $\sigma$ along the zigzag direction, at three distinct temperatures $T$ shown by different colors. An increase in temperature leads to an increase in the amplitude of strain oscillations as well as in an increase in the average strain. The latter one is obtained as the average over both realizations and time, $\overline{\left\langle{\epsilon_{T}}\right\rangle}$ and it is indicated by the dashed horizontal lines of different colors depending on the temperature in Figure 3. The average strain $\overline{\left\langle{\epsilon_{700}}\right\rangle}$ (red horizontal dashed lines) about which the $T=700$ K curves oscillate are higher than the $\overline{\left\langle{\epsilon_{100}}\right\rangle}$ (blue horizontal dashed lines) of the $T=100$ K curves in all cases of different stress. However, these differences are larger on absolute values for larger stresses.
@@
Figure 3: Time evolution of the average (over individual realizations) strain $\left\langle\epsilon_{T}(t)\right\rangle$, Equation (6), when a stress (a) $\sigma=0.188$ eV/Å2, (b) $\sigma=1.03$ eV/Å2, and (c) $\sigma=1.97$ eV/Å2, along the zigzag direction is applied, for different temperatures $T=100$ K (blue curves), $T=400$ K (green curves), and $T=700$ K (red curves). The average (over realizations and time) strain $\overline{\left\langle\epsilon_{T}\right\rangle}$, for each temperature, is indicated by the horizontal dashed lines of the same color in each panel.
@@
Calculating the average strain $\overline{\left\langle{\epsilon_{T}}\right\rangle}$ as mentioned above, the mechanical response of planar graphene at different temperatures is obtained. Stress-strain curves for uniaxial tensile loads along the armchair and zigzag directions are presented for various temperatures in Figure 4. Despite the small differences, one can see for larger stresses that the average strain is a bit further to the right for the higher temperature cases. The error-bars indicate the standard deviation of the average strain measurement. As one can also deduce from Figure 3, the standard deviation is higher for higher temperatures. We further note that these strain measurements have been tested for larger numbers of realizations or longer time windows. In particular, the measured strain differs in the 3rd significant digit at most, when increasing the number of realizations or doubling the length of the time window.
@@
Figure 4: Stress-strain response of planar graphene for uniaxial loads along the (a) armchair (b) zigzag direction, for different temperatures as indicated in the legend. Filled circles indicate the measured average strain for each given stress. Solid curves represent fittings of these data points with Equation (7). For $T\neq 0$ K the strain is given as the average over time and realizations, $\overline{\left\langle\epsilon_{T}\right\rangle}$, and the error-bars correspond to one standard deviation.
@@
Since in Figure 4 the strain is measured with respect to the average length $\ell_{T}^{r}$ due to thermally induced oscillations, the stress-strain curves pass from the origin of Figure 4, as expected. From Figure 4, we see that the temperature has a relatively small effect in the stress-strain response, at least for the values of $T$ considered here, apart from the significant reduction of the fracture point. For small stresses the achieved strain is practically the same for the two directions of applied stress, while the strong directional dependence at large stresses has already been well established in previous investigations ref-Liu-PRB-2007 ; ref-Zhao-NanoLet-2009 ; ref-Kalosakas-JAP-2013 ; ref-deSousa-Intech-2021 .
The stress-strain response can be described by the nonlinear relation
@@
where $\sigma$ is the applied uniaxial stress, $\epsilon$ the corresponding strain, $E_{2D}$ is the 2D Young’s modulus and $D_{2D}$ is the 2D third-order elastic modulus. For each temperature examined and both directions of applied stress, we fit the data presented in Figure 4 with Equation (7) to determine graphene’s elastic moduli. The computed values of $E_{2D}$ and $D_{2D}$ are plotted in Figure 5 as a function of temperature, for applied stress in either the armchair (red points) or the zigzag (blue points) direction. The error-bars on these points indicate one standard deviation of the fitted parameters under the observed covariance of the fit. Furthermore, linear fittings of these data points are indicated by the dashed red (dotted blue) line for stress in the armchair (zigzag) direction.
@@
Figure 5: Temperature dependence of (a) the Young’s modulus $E_{2D}$, and (b) the third-order elastic modulus $D_{2D}$, for applied stress along the armchair (red points) or the zigzag (blue points) direction, evaluated through fittings of the data of Figure 4 with Equation (7).
@@
Figure 5 shows that both elastic moduli exhibit an almost linear dependence on temperature. The Young’s modulus $E_{2D}$ [Figure 5(a)] appears to decrease linearly with increasing temperature, albeit only by a relatively small amount, which is consistent with other results in the literature ref-Shen-MatDes-2010 ; ref-Shao-JCP-2012 ; ref-Tahani-AMM-2024 . In particular, the linear fitting of these data for stress along the armchair direction leads to a variation of $E_{2D}(T)$ with a slope $-8.1\times 10^{-3}$ (N/m)/K, while for stress along the zigzag direction the slope is $-4.37\times 10^{-3}$ (N/m)/K. The decrease of Young’s modulus with temperature is often given in the literature as a percentage change over the investigated range of temperatures. In our case the Young’s modulus decreases by $1.8\%$ ($0.94\%$) for stress along the armchair (zigzag) direction over the temperature range from 0 K to 700 K. In ref-Shen-MatDes-2010 , MD is used for investigations over the range from 300 K to 700 K, and $E_{2D}$ is found to decrease by 5.1%, 1.9%, or 1.3% (decrease by 3.5%, 2.6%, or 1.3%) for strain in the armchair (zigzag) direction, where the three different measurements are for three different aspect ratios of the graphene lattice. In ref-Shao-JCP-2012 density functional theory is used to determine that $E_{2D}$ decreases by 2.2% over the range from 0 K to 1000 K, with no directional dependence indicated. In ref-Tahani-AMM-2024 this reduction in $E_{2D}$ is found to be between 3% and 4% (no directional dependence indicated) for the temperature ranging from 0 K to 1600 K, where the varying reduction depends on the exact assumptions made for the asymptotic homogenization model, which affects the value at $T=0$ K.
@@
We observe from Figure 5(b) that the $D_{2D}$ values are consistently higher for strain in the zigzag direction, than for the other direction. This is congruent with the fact that the graphene sheet is more resistant to stress along the zigzag direction. Since there are no bonds parallel to the direction of strain when forces are applied along the zigzag direction, the sheet can tolerate larger stresses. When stress is along the armchair direction, one third of all the bonds are parallel to the direction of strain, and hence these bonds exhibit maximal stretching in the sheet, leading to higher strain for the same stress, and hence the lower $D_{2D}$ values. This higher strain per stress due to the particular direction can be seen when comparing the panels of Figure 4, where the curves for stress in the armchair direction lie further to the right than when the stress is applied in the zigzag direction, i.e., indicating higher strains for similar stresses. Bonds that are not parallel to the direction of strain also stretch out, but due to their orientation, more stress is required to stretch them as significantly as the parallel bonds are stretched. Regarding the temperature variation of $D_{2D}$, different trends are exhibited when the stress is along the zigzag or the armchair direction. A linear fitting of the $D_{2D}(T)$ data points results in a slope $+4.95\times 10^{-2}$ (N/m)/K for strain in the armchair direction and $-1.01\times 10^{-2}$ (N/m)/K for strain in the zigzag direction. The value of $D_{2D}$ increases by $6.3\%$ (decreases by $1.4\%$) for strain in the armchair (zigzag) direction over the temperature range from 0 K to 700 K.
@@
Furthermore, we estimate the graphene’s tensile strength $\sigma_{f}$ and failure strain $\epsilon_{f}$, for different temperatures $T$. The former one is obtained by the highest tested value of $\sigma$ which does not lead to failure of the graphene sheet. Its error-bar is provided by the step we use in the increment of the tested $\sigma$ values, which are evenly spaced. These results are presented in Figure 6(a), where an almost linear decrease of the fracture stress with temperature is shown. A linear fitting of these data points is indicated with a dashed red (dotted blue) line for stress in the armchair (zigzag) direction. The slope of the linear fitting of the $\sigma_{f}(T)$ data is $-8.37\times 10^{-3}$ (N/m)/K for stress in the armchair direction and $-1.45\times 10^{-2}$ (N/m)/K for stress in the zigzag direction. Such a linear dependence of the tensile strength on temperature is in accordance with existing results ref-Zhao-JAP-2010 . In particular, we can estimate that the reported tensile strength per temperature has a slope of $-8.57\times 10^{-3}$ (N/m)/K in this aforementioned MD investigation of a 3936 atom graphene lattice.
@@
Figure 6: Temperature dependence of (a) the tensile strength $\sigma_{f}$ and (b) the associated failure strain $\epsilon_{f}$ of graphene. Straight lines represent linear fittings.
@@
We have also estimated the failure strain, $\epsilon_{f}$, at different temperatures, by solving for $\epsilon_{f}$ the Equation (7). To this end, the known value of $\sigma_{f}$ as well as the fitted values of $E_{2D}$ and $D_{2D}$ describing the stress-strain curve at the given temperature, have been used. In this case the error-bars are determined by converting the corresponding extreme values of stress, $\sigma_{f}\pm\Delta\sigma_{f}$, to strain [via Equation (7)], and then choosing the maximum absolute difference from $\epsilon_{f}$. These results are shown in Figure 6(b) and again a linear fitting of the data is indicated with a dashed red (dotted blue) line for stress in the armchair (zigzag) direction. The linear fitting of the $\epsilon_{f}(T)$ data points leads to a slope $-4.71\times 10^{-3}$ strain/K for stress in the armchair direction and $-1.60\times 10^{-2}$ strain/K for stress in the zigzag direction. It is reasonable that higher temperatures cause failure at lower stresses/strains, since the maximum deformation of the graphene lattice under a fixed load increases with temperature, as shown in Figure 3. Therefore the sheet breaks more easily at higher temperatures. Moreover, the known property that graphene can tolerate higher loads along the zigzag direction is clear by the results in Figure 6, where the values of $\sigma_{f}$ and subsequently $\epsilon_{f}$ are consistently lower for stress in the armchair direction (red data points) as compared to stress applied in the zigzag direction (blue data points).

### v4

@@ 3.1 Temperature calibration
When an energy density $e_{N}$ is inserted in the strained graphene lattice, we observe that initially the total kinetic energy increases with time from its zero starting value and then following some relatively large fluctuations the system is settled at thermal equilibrium after at most $10^{3}~t_{u}$. The time evolution of the system’s temperature $T(t)$ towards equilibrium is computed in our microcanonical MD simulations through the energy equipartition relation
@@
In order to test whether thermal equilibrium has been reached, we compare the mean value and the standard deviation of the fluctuating temperature over various time windows. Before achieving thermal equilibrium the standard deviation of the time-averaged $T$ is relatively large and also changes depending on the time window. When thermal equilibrium is reached the temperature fluctuations and the standard deviation are consistently small. The mean temperature at thermal equilibrium is calculated by averaging over both the individual realizations and the recording time-window. We denote this average value $\overline{\left\langle T\right\rangle}$ by $T_{ave}$. In this case, the standard deviation of the measured values is computed using all data points over realizations and time.
The relationship between the additional energy density $e_{N}$ on top of the relaxed equilibrium loaded structures and the averaged temperature $T_{ave}$ is linear in all cases of different stresses examined here, at least for temperatures up to 700 K considered in this work. One representative case is shown in Figure 2. The resulting slopes from the linear fittings of the data are very close for all values of stress $\sigma$ (a difference in the computed values is observed only in the fourth significant digit) and they are slightly above $2k_{B}$ due to the nonlinearities of the potential energy. For finite loads, the slope slightly increases with the amount of stress. Thus, for a given value of stress $\sigma$, we use the obtained slope of the $e_{N}$ versus $T_{ave}$ linear fitting in order to control temperature (within a 1% accuracy) in our investigation. In particular, we are setting the amount of the added energy density $e_{N}$ according to the desired temperature value.
Figure 2: Symbols represent the relation between the additional energy density $e_{N}$ above the relaxed $T=0$ K graphene structure subjected to uniaxial tensile stress $\sigma=2.16$ eV/Å2 along the zigzag direction, and the average temperature $T_{ave}$ at thermal equilibrium, evaluated through the MD simulations by averaging over both time and the different realizations. One standard deviation of the $T_{ave}$ measurements is indicated by blue horizontal error-bars. The linear fitting of the presented data points is shown by gray solid line, providing a slope equal to $1.74\times 10^{-4}$ eV/K.
In order to investigate the system’s elastic and structural properties, discussed in the following sections, we collect data from the central region of the lattice for avoiding potential edge effects and thus representing the behavior of bulk graphene. In particular, this central region sub-lattice has an analogous geometry to the larger structure, with number of columns $n_{I}=44$ and number of rows $n_{J}=45$.
@@ 3.2 Mechanical response
In the stress-controlled numerical implementation used here to examine the mechanical response of graphene, we compute the resulting strain due to the fixed force applied at the appropriate edge atoms. The uniaxial strain $\epsilon$ is obtained through the strain of the central row of the graphene sheet when the stress is applied along the armchair direction, while it is calculated by the average strain on the two central columns of the sheet in case of stress along the zigzag direction (see Figure 1). By $\epsilon_{T}$ is indicated the strain corresponding to temperature $T$.
@@
In the zero temperature case, $T=0$ K, the uniaxial strain $\epsilon_{0}$ is determined through the relative change of the length of the central row (two central columns) of graphene subjected to a given applied stress along the armchair (zigzag) direction, with respect to the length of the central row (columns) at the unstrained equilibrium configuration shown in Figure 1. These measurements are taken in the central region of the lattice, as mentioned at the end of previous subsection. For any length computations discussed here, a horizontal (vertical) length is measured as the difference of the $x$ ($y$) coordinates of the considered edge atoms. The stress-strain response is obtained in this way at 0 K
@@
When the temperature of the system is raised at finite values, by adding energy to the equilibrated graphene, the lattice stretches and compresses in an oscillatory manner. The details of these oscillations depend on the temperature and the applied stress, and will be investigated in the future. In this case one has to take into account that the strain measurement $\epsilon_{T}(t)$ is now exhibiting temporal oscillations. Since we consider 10-20 different realizations of the randomly inserted initial energy distribution, we register the average, over these realizations, strain in time $\left\langle\epsilon_{T}(t)\right\rangle$, noting that the aforementioned oscillations are in-phase in the different realizations.
For evaluating the strain $\epsilon_{T}$ of a uniaxially loaded graphene sheet at finite temperatures, a reference length $\ell_{T}^{ref}$ corresponding to zero stress $\sigma=0$ at the particular value of $T$ is needed. This reference length accounts for thermal effects on the initial configuration and it is obtained by calculating the average, over realizations and time, of the length of the central row, or columns, of the sheet in the absence of any load. Then, when a stress is applied the time evolution of strain in a particular realization is determined as the relative change of the length with respect to the reference length $\ell_{T}^{ref}$
$\epsilon_{T}(t)=\frac{\ell(t)-{\ell_{T}^{ref}}}{{\ell_{T}^{ref}}},$ (6)
@@
where $\ell(t)$ is the length at time $t$ of the central row or the average length of the two central columns depending on the direction of the applied uniaxial load.
@@
In Figure 3 we highlight the behavior of $\left\langle{\epsilon_{T}}(t)\right\rangle$ for various values of stress $\sigma$ along the zigzag direction, at three distinct temperatures $T$ shown by different colors. An increase in temperature leads to an increase in the amplitude of strain oscillations as well as in an increase in the average strain. The latter one is obtained as the average over both realizations and time, $\overline{\left\langle{\epsilon_{T}}\right\rangle}$ and it is indicated by the dashed horizontal lines of different colors depending on the temperature in Figure 3. The average strains $\overline{\left\langle{\epsilon_{700}}\right\rangle}$ (red horizontal dashed lines) about which the $T=700$ K curves oscillate are higher than the $\overline{\left\langle{\epsilon_{100}}\right\rangle}$ (blue horizontal dashed lines) of the $T=100$ K curves in all cases of different stress. However, these differences are larger on absolute values for larger stresses.
@@
Figure 3: Time evolution of the average (over individual realizations) strain $\left\langle\epsilon_{T}(t)\right\rangle$, Equation (6), when a stress (a) $\sigma=0.188$ eV/Å2, (b) $\sigma=1.03$ eV/Å2, and (c) $\sigma=1.97$ eV/Å2, along the zigzag direction is applied, for different temperatures: $T=100$ K (blue curves), $T=400$ K (green curves), and $T=700$ K (red curves). The average (over realizations and time) strains $\overline{\left\langle\epsilon_{T}\right\rangle}$, for each temperature, are indicated by the horizontal dashed lines of the same color in each panel.
We have checked that if one follows an alternative path on the $(\sigma,T)$-plane by giving first initial energy to the system and then applying forces at the edges, practically identical average strains are obtained. However, our approach is much more efficient because the temperature is accurately controlled and, more importantly, the system reaches thermal equilibrium significantly faster; in the alternative method the equilibration takes orders of magnitude longer.
@@
Calculating the average strain $\overline{\left\langle{\epsilon_{T}}\right\rangle}$ as mentioned above, the mechanical response of planar graphene at different temperatures is obtained. Stress-strain curves for uniaxial tensile loads along the armchair and zigzag directions are presented for various temperatures in Figure 4. Despite the small differences, one can see for larger stresses that the average strain is a bit further to the right for the higher temperature cases. The error-bars indicate the standard deviation of the average strain measurement. As one can also deduce from Figure 3, the standard deviation is higher for higher temperatures. This is highlighted via the insets in each panel, where a close-up of the data points and error-bars is presented for the region which is indicated by the grey rectangle in each panel. Close-up it is easier to see that the lengths of the error-bars increase with temperature.
We have checked the accuracy of the presented strain measurements when more realizations or longer time windows are considered. In particular, the obtained strain values differ in the 3rd significant digit at most, when increasing the number of realizations or doubling the length of the time window.
@@
Figure 4: Stress-strain response of planar graphene for uniaxial loads along the (a) armchair (b) zigzag direction, for different temperatures as indicated in the legend. Filled circles indicate the obtained average strain for each given stress. Solid curves represent fittings of these data with Equation (7), see text. For $T\neq 0$ K the strain is given as the average over time and realizations, $\overline{\left\langle\epsilon_{T}\right\rangle}$, and the error-bars correspond to one standard deviation. The insets in each panel depict a close-up view of the region indicated by the grey rectangle in each panel.
@@
Since at finite temperatures the strain is measured with respect to the averaged oscillating length $\ell_{T}^{ref}$ due to thermally induced vibrations of the unstrained sheet, the stress-strain curves pass from the origin of Figure 4, as expected. From Figure 4, we see that the temperature has a relatively small effect on the stress-strain response, at least for the values of $T$ considered here, apart from the significant reduction of the fracture point. For small stresses the achieved strain is practically the same for the two directions of applied stress, while the strong directional dependence at large stresses has already been well established in previous investigations ref-Liu-PRB-2007 ; ref-Zhao-NanoLet-2009 ; ref-Kalosakas-JAP-2013 ; ref-deSousa-Intech-2021 .
The stress-strain response can be described by the nonlinear relation ref-Kalosakas-JAP-2013
@@
where $\sigma$ is the applied uniaxial stress, $\epsilon$ the corresponding strain, $E_{2D}$ is the 2D Young modulus and $D_{2D}$ is the 2D third-order elastic modulus. For each temperature examined and both directions of applied stress, we first obtained the value of Young modulus by the linear response at small stress/strain and then we fit the data presented in Figure 4 with Equation (7) to determine the third-order elastic modulus. The computed values of $E_{2D}$ and $D_{2D}$ are plotted in Figure 5 as a function of temperature, for applied stress in either the armchair (red points) or the zigzag (blue points) direction. The error-bars on these points indicate one standard deviation of the fitted parameters under the observed covariance of the fit. A linear variation can roughly approximate the obtained temperature dependence of these elastic moduli. Linear fittings of the corresponding data are indicated by the dashed red (dotted blue) line for stress in the armchair (zigzag) direction.
@@
Figure 5: Temperature dependence of (a) the Young modulus $E_{2D}$ and (b) the third-order elastic modulus $D_{2D}$, for applied stress along the armchair (red points) or the zigzag (blue points) direction, evaluated through fittings of the data of Figure 4 with Equation (7) (see text).
@@
The Young modulus $E_{2D}$ [Figure 5(a)] appears to decrease almost linearly with increasing temperature, albeit only by a relatively small amount, which is consistent with other results in the literature ref-Shao-JCP-2012 ; ref-Shen-MatDes-2010 ; ref-Tahani-AMM-2024 . In particular, the linear fitting of these data for stress along the armchair direction leads to a variation of $E_{2D}(T)$ with a slope $8.1\times 10^{-3}$ (N/m)/K, while for stress along the zigzag direction the slope is $4.4\times 10^{-3}$ (N/m)/K.
The decrease of Young modulus with temperature is often given in the literature as a percentage change over the investigated range of temperatures. To enhance a direct quantitative comparison, we present our results along with existing ones in the literature as percentage change per 100 K. In our case the Young modulus decreases by $0.25\%/100$ K and $0.13\%/100$ K for stress along the armchair or the zigzag direction, respectively. Combined density functional theory and quasi-harmonic approximation calculations in Reference ref-Shao-JCP-2012 have determined a $E_{2D}$ decrease by $0.22\%/100$ K over the range from 0 K to 1000 K. Molecular dynamics has been used in Reference ref-Shen-MatDes-2010 for investigations at 300 K, 500 K, and 700 K of graphene lattices of different aspect ratios consisting of 1886 atoms. $E_{2D}$ is found to decrease by $1.3\%/100$ K, $0.48\%/100$ K, and $0.33\%/100$ K (decrease by $0.88\%/100$ K, $0.65\%/100$ K, and $0.33\%/100$ K) for strain along the armchair (zigzag) direction, where the three different values correspond to graphene aspect ratios 1.97, 1.44, and 1.01 (1.95, 1.45, and 0.99). We note that our lattice has an aspect ratio 1.72. Molecular dynamics simulations in Reference ref-Zhang-CompMatSci-2013 have resulted in a $E_{2D}$ decrease by $1.4\%/100$ K over a temperature range of 300 K to 2000 K, for strain applied along the armchair direction. Finally, in Reference ref-Tahani-AMM-2024 a reduction in $E_{2D}$ between $0.19\%/100$ K and $0.25\%/100$ K has been obtained for temperatures ranging from 0 K to 1600 K, where the varying reduction depends on the different parameterizations of the used model, which affects the Young modulus value at $T=0$ K.
The calculated Young modulus $E_{2D}$ at 300 K is in good agreement with values reported in experimental studies conducted at room temperature ref-Lee-Science-2008 ; ref-Cao-NatComm-2020 ; ref-Varillas-MechSci-2024 . In particular, our findings, namely 315315~N/m and 312312~N/m for loading along the zigzag and armchair direction, respectively, are consistent with the Young modulus reported in References ref-Lee-Science-2008 ($340\pm 50$~N/m), ref-Cao-NatComm-2020 ($300$ to 340340~N/m), and ref-Varillas-MechSci-2024 ($350\pm 100$~N/m).
@@
We observe from Figure 5(b) that the $D_{2D}$ values are consistently higher for strain in the zigzag direction, than for the other direction. This is congruent with the fact that the graphene sheet is more resistant to stress along the zigzag direction. When stress is along the armchair direction, one third of all the bonds are parallel to the direction of strain, and hence these bonds exhibit maximal stretching in the sheet. Taking into account the respecting angle deformations there is in general a higher strain for the same stress in this case as compared to loads along the zigzag direction. Considering that the Young modulus is almost the same in this two cases, this leads to lower $D_{2D}$ modulus (i.e. higher absolute values) for stresses in the armchair direction. The different strains for a given stress in the two perpendicular loading directions discussed here can be seen when comparing the panels of Figure 4, where the curves for stress in the armchair direction lie further to the right than when the stress is applied in the zigzag direction, i.e., indicating higher strains for the same stress. Regarding the temperature variation of $D_{2D}$, different trends are exhibited when the stress is along the zigzag or the armchair direction. A linear fitting of the $D_{2D}(T)$ data points results in a slope $5.0\times 10^{-2}$ (N/m)/K for strain in the armchair direction and $1.0\times 10^{-2}$ (N/m)/K for strain in the zigzag direction. The value of $D_{2D}$ increases by $0.91\%/100$ K (decreases by $0.19\%/100$ K) for strain in the armchair (zigzag) direction over the temperature range from 0 K to 700 K.
@@
Finally, we estimate the graphene’s fracture strength $\sigma_{f}$ and failure strain $\epsilon_{f}$, for different temperatures $T$. The former one is obtained by the highest tested value of stress $\sigma$ which does not lead to failure of the graphene sheet. Its error-bar is provided by the step we use in the increment of the tested $\sigma$ values, which are evenly spaced. These results are presented in Figure 6(a), where an almost linear decrease of the fracture stress with temperature is shown. A linear fitting of these data points is indicated with a dashed red (dotted blue) line for stress in the armchair (zigzag) direction. The slope of the linear fitting of the $\sigma_{f}(T)$ data is $8.4\times 10^{-3}$ (N/m)/K for stress in the armchair direction and $1.5\times 10^{-2}$ (N/m)/K for stress in the zigzag direction.
Such a linear dependence of the fracture strength on temperature is in accordance with existing results. In particular, we estimate that, in the MD study of graphene loaded along the armchair direction reported in Reference ref-Zhao-JAP-2010 , the fracture strength decreases with temperature at a rate of $8.6\times 10^{-3}$ (N/m)/K, which is in very good agreement with the slope obtained in our work for the same direction. Molecular dynamics simulations are also used in Reference ref-Tanhadoust-DiamRelMat-2022 and the presented temperature dependence of the fracture stress decreases linearly with a slope of $1.4\times 10^{-2}$ (N/m)/K, however it is not clear which was the loading direction there. In another MD investigation, where the stated loading direction was the armchair direction, the corresponding slope was found $1.6\times 10^{-2}$ (N/m)/K ref-Zhang-CompMatSci-2013 . Lastly, Reference ref-Varma-IntJMechMatDes-2022 employs a combination of machine learning and MD simulations with the Tersoff potential. In that study, the fracture strength decreases linearly with temperature, exhibiting a slope of $3.8\times 10^{-2}$ (N/m)/K for loading along the zigzag direction, whereas for loading along the armchair direction a bilinear behavior is observed: the slope is initially $6.7\times 10^{-3}$ (N/m)/K but becomes steeper above 500 K, thus representing the only non-linear trend reported in the literature.
@@
Figure 6: Temperature dependence of (a) the fracture strength $\sigma_{f}$ and (b) the associated failure strain $\epsilon_{f}$ of graphene. Straight lines represent linear fittings.
@@
The failure strain, $\epsilon_{f}$, at different temperatures has been estimated through the value of fracture stress by solving for $\epsilon_{f}$ in Equation (7). To this end, the known value of $\sigma_{f}$ as well as the fitted values of $E_{2D}$ and $D_{2D}$ describing the stress-strain curve at the given temperature, have been used. In this case the error-bars are determined by converting the corresponding extreme values of stress, $\sigma_{f}\pm\Delta\sigma_{f}$, to strain [via Equation (7)], and then choosing the maximum absolute difference from $\epsilon_{f}$. These results are shown in Figure 6(b), where again a linear fitting of the data is indicated with a dashed red (dotted blue) line for stress in the armchair (zigzag) direction. The linear fitting of the $\epsilon_{f}(T)$ data points leads to a slope $4.7\times 10^{-3}$ % strain/K for stress along the armchair direction and $1.6\times 10^{-2}$ % strain/K for stress in the zigzag direction. We can further compare these slopes to existing results by estimating the slopes of the failure strain versus temperature data reported in the literature. For uniaxial loading along the armchair direction in Reference ref-Zhao-JAP-2010 , a slope of $4.8\times 10^{-3}$ % strain/K is estimated, which is in very good agreement with our result. A slope of $5.4\times 10^{-3}$ % strain/K is estimated for the results presented in Reference ref-Tanhadoust-DiamRelMat-2022 , where, as noted above, it is not clear which is the loading direction. Finally, the corresponding slope for strain applied along the armchair direction is $5.9\times 10^{-3}$ % strain/K in Reference ref-Zhang-CompMatSci-2013 .
The fracture stress reported in the experimental study of Reference ref-Lee-Science-2008 is $42\pm 4$ N/m. We have obtained an intrinsic strength of $39.1\pm 1.5$~ N/m ($28.7\pm 1.3$ N/m) and a fracture strain of $18.5\pm 1.5\%$ ($12.0\pm 0.8\%$) for loading along the zigzag (armchair) direction at 300 K. We note that our results are in good agreement with the intrinsic strength and fracture strain values reported in other studies which additionally allow out-of-plane deformations of the material. Specifically, MD simulations carried out at 300 K using the AIREBO potential, yield, for loading along the zigzag direction, intrinsic strengths around 363736-37 N/m and fracture strains between $17-20\%$ ref-Zhao-NanoLet-2009 ; ref-Varillas-MechSci-2024 . For loading along the armchair direction, Reference ref-Lee-NanoResLets-2015 conducted MD simulations using the REBO force field at 300 K, obtaining a fracture strain of $12.5\%$ at a stress of $29.1$ N/m, while Reference ref-Zhao-NanoLet-2009 using AIREBO reports an intrinsic strength of 3030~N/m and fracture strain $13\%$.
The results shown in Figure 6 indicate that graphene fails at lower applied stress/strain as temperature increases. This is reasonable since, as can be clearly seen from Figure 3, for the stress-controlled simulations considered here the sheet achieves higher strains over the course of its oscillations by increasing temperature. As a result, for fixed stress the bonds between neighboring atoms experience longer stretches at higher temperatures and therefore are more likely to break, causing failure of the material, due to the increase of the maximum deformation of the lattice. Moreover, graphene can tolerate higher loads along the zigzag direction in the whole temperature range investigated here, as is implied by the results of Figure 6, where the values of fracture strength and failure strain are consistently lower for stress in the armchair direction (red data points) as compared to stress applied in the zigzag direction (blue data points). In contrast to the relatively stronger temperature dependence of the fracture strength and failure strain, the Young modulus variation shown in Figure 5(a) exhibits a much smaller relative change, implying that the influence of thermal effects on the stiffness of graphene is less significant, at least within this temperature regime.